OpenStudy (anonymous):

A baseball team plays in he stadium that holds 72000 spectators. With the ticket price at 9 dollars, the average attendence has been 30000. When the price dropped to 7 dollars, the average attendence rose to 36000. a) Find the demand function p(x), where x is the number of the spectators. (assume that p(x) is linear) p(x)= b) What should be the ticket price so that the revenue is maximized?

OpenStudy (anonymous):

i got p(x) to be (-x/300)+129 and its not correct...any suggestions?

OpenStudy (blockcolder):

If p(x) is linear and two points that satisfy this is (30000,9) and (36000,7), then we have: \[p(x)=\frac{9-7}{30000-36000}(x-30000)+9 \\ \Rightarrow p(x)=-\frac{1}{3000}(x-30,000)+9\\ \Rightarrow p(x)=-\frac{x}{3000}+19\] If we let R(x):=revenue function, then \[R(x)=xp(x)=-\frac{x^2}{3000}+19x\] Can you find the x where this is maximum and the corresponding p(x)?

OpenStudy (anonymous):

how would i do this?

OpenStudy (anonymous):

Take the derivative wrt x and set to zero.

OpenStudy (anonymous):

ok so then x= 28500

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

but thats incorrect according to my hw

OpenStudy (blockcolder):

And when x=28500, p(x)=???

OpenStudy (anonymous):

Your answer book is asking for p(x). :)

OpenStudy (anonymous):

And when x=28500, p(x)=2707519

OpenStudy (anonymous):

NOOOO

OpenStudy (anonymous):

-28500/3000 + 19 = ?

OpenStudy (anonymous):

@GT 29.4

OpenStudy (anonymous):

You calculated R(x) above to get 2707519?

OpenStudy (anonymous):

-28500/3000 + 19 = -9.5 + 19 = $9.5

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Remember, in that equation for p(x), the value 19 signifies where attendance drops to ZERO. So, no price above $19 even makes sense.

OpenStudy (anonymous):

ok i got it thanks

OpenStudy (anonymous):

The manager of a large apartment complex knows from experience that 80 units will be occupied if the rent is 320 dollars per month. A market survey suggests that, on the average, one additional unit will remain vacant for each 8 dollar increase in rent. Similarly, one additional unit will be occupied for each 8 dollar decrease in rent. What rent should the manager charge to maximize revenue?

OpenStudy (anonymous):

Good!

OpenStudy (anonymous):

how abt this?

OpenStudy (anonymous):

Post it separately. It is the policy of this site.

OpenStudy (anonymous):

o ok